# ASTRODEEP-JWST: NIRCam-HST multiband photometry and redshifts for half a million sources in six extragalactic deep fields

### v1

@@ 2.4 PRIMER-COSMOS and PRIMER-UDS
Data is from the PRIMER program (GO 1837, P.I. Dunlop; no F775W for PRIMER-COSMOS and no F775W and F105W for PRIMER-UDS). The COSMOS FoV has an area of $\simeq$141.8 sq. arcmin, and the UDS FoV has an area of $\simeq$251.2 sq. arcmin. The images have been reduced by D. Magee, with HST data re-reduced from the CANDELS imaging.
@@ 3 Methods
In this Section we summarize the techniques and algorithms we used to prepare the images and extract the photometric information.
@@ 3.1 Alignment and astrometry
We assessed the quality of the astrometric registration by cross-matching the source coordinates ALPHAWIN_J2000 and DELTAWIN_J2000 obtained running SExtractor (Bertin & Arnouts, 1996) on all the bands in each field before re-projecting the images to the same common grid. In most cases we found small $\Delta$RA and $\Delta$DEC offsets (of the order of a few mas) between the native astrometry of the JWST and the HST bands. However, since in some cases the offsets were larger and not negligible, we applied a custom algorithm to correct all of them. First, we computed the offset between the HST WFC3 F160W band and the JWST NIRCam F150W band, and re-aligned the F160W image. Then we corrected the offsets of the other HST bands by cross-matching the coordinates of the sources with those extracted from the re-aligned F160W band (see Figure 1 and Tab. 6 in Appendix). Finally, we re-projected all the images on the same NIRCam pixel grid.
Figure 1: Example of the validation tests made to fix the astrometric registration of the HST bands. Shown is the displacement $\Delta$RA and $\Delta$DEC between the sources detected in the CEERS F125W image and those detected in the CANDELS EGS F160W image, re-aligned to the reference frame of the CEERS F150W image, before (left panel) and after (right panel) applying the correction described in Section 3.1.
@@ 3.2 RMS scaling
Because the photometric errors are computed by means of the RMS maps (see Sect. 3.4), we checked that the latter were indeed representative of the true uncertainties of the measurement, using an improved version of the technique described in M22. In short, the RMS image of each band was subdivided into 2 to 4 complementary sub-regions of comparable exposure time, by means of an automatic algorithm applied to the weight maps. Then, in each of these sub-region 300 artificial point sources (WebbPSF simulated point spread functions) were injected at random positions, excluding areas assigned to real sources by means of SExtractor segmentation maps, and the dispersion of their fluxes (measured within an aperture of 0.1” using the software a-phot, Merlin et al., 2019) was compared to their nominal errors to obtain a re-scaling factor for that region of the RMS map. Finally, the original map – which includes Poissonian photon noise – was then re-scaled by the median value of such factors. The typical values for SW bands are below $\sim$1.5, while they can get as high as $\sim$2.0 for some LW bands and fields.
@@ 3.3 Detection
Because the catalogs are mainly designed for the study of the high redshift universe, major effort was put in optimizing the measurements of faint extended objects, rather that those of bright local galaxies or stars. The strategy is similar to that adopted in M22 and P23: we chose to detect in the infrared, smoothing the scientific image with a Gaussian convolution filter with FWHM = 0.14” (close to the F356W and F [^4] W FWHMs, see below), and applying a detection threshold corresponding to SNRsimilar-$\sim$2, to pick up as many faint high redshift sources as possible. However, there is a substantial difference with respect to previous work: rather than just using F [^4] W as detection band, we create a weighted stack of F356W and F [^4] W, allowing to single out sources that peak at 3.5 to 4 $\mu$m, and exploiting the F356W mosaics which are often as deep as, or deeper than, the F [^4] W ones (see Appendix D). The resulting depths of the detection stacks are shown in Fig. 2.
@@
We then run SExtractor v2.8.6, in the customised version used for the CANDELS campaign. Most of the parameters were left to the default values, with the exception of those given in Table 3. We tried to use the same values for all fields, but we found that DETECT_THRESH and ANALYSIS_THRESH needed adjustment in some cases, to ensure optimal results. In particular, we set them to 0.85 for PRIMER-COSMOS, and to 0.55 for ABELL2744; these values yielded the best trade-off between purity and completeness, allowing for the detection of faint sources while avoiding too many spurious objects to be included in the list.
@@
The final detection catalogs contain a grand total of 531,173 sources, as reported in Table 4. In the catalogs we include a unique object identifier number, the equatorial position (Right Ascension and declination, in degrees), and basic morphological information such as the area of the segmented cluster of pixels and the half light radius from SExtractor, and the semimajor axis, ellipticity and position angle of the elliptical isophote obtained by means of a-phot; see Appendix A for more details.
@@
Fig. 3 shows the point-source detection completeness for the six fields, determined by injecting fake PSF-shaped objects in empty regions of the F356W+F [^4] W stack (using the detection segmentation map as a mask), running SExtractor with the same parameters used in the actual detection process, and checking the fraction of them being actually detected. The different depths of the detection images is evident, NGDEEP being the deepest with 90% at ABsimilar-$\simeq$30 and 50% at ABsimilar-$\simeq$30.5, and PRIMER-UDS the shallowest with 90% at ABsimilar-$\simeq$28.5 and $50$% at ABsimilar-$\simeq$29; noticeable is the complicated pattern of the JADES-GS field, which comprises deep and shallow regions (90% completeness at ABsimilar-$\simeq$28.5 but 50% at ABsimilar-$\simeq$30).
@@
Figure 3: Detection completeness. Shown is the fraction of fake point sources of given magnitude injected in the detection F356W+F [^4] W stacks, and detected with SExtractor runs performed with the parameters used in real runs.
@@ 3.4 Photometry
As in our previous efforts, in our public catalogs we provide total fluxes. The main methods used for this work are the same ones adopted in M22 and P23. In short, we compute colors from fixed circular apertures on PSF-matched images, and assuming no color gradients outside them we then re-scale to total fluxes multiplying them by the flux within Kron (1980) elliptical apertures in the F356W+F [^4] W stack, as measured by the software a-phot (Merlin et al., 2019). So, the total flux in each band is given by the formula $f_{tot,band}=c_{ap,det}\times f_{ap,band}$, where $c_{ap,det}=f_{tot,det}/f_{ap,det}$. As pointed out in M22, our a-phot Kron-like aperture on average tends to gather more light than the standard SExtractor MAG_AUTO (see Fig. 4 in M22), so we don’t apply any further aperture correction to the total fluxes. Errors are estimated summing in quadrature the relevant pixels from the RMS maps and applying the same formula, $e_{tot,band}=c_{ap,det}\times e_{ap,band}$; this choice is motivated by the fact that most scientific applications (e.g. SED-fitting or color-selections) are essentially based on colors, so the propagation of the total flux error would overestimate the relevant uncertainty. We compute fluxes within nine apertures (with diameters 0.2”, 0.28”, 0.33”, 0.50”, 0.66”, 0.70”, 1.32”, 2.65” and 5.30”), and build just as many catalogs of total fluxes. The values of the apertures correspond to integer multiples of the F [^4] W FWHM 0.165”, except for the two smallest ones (a fixed 0.1” radius aperture and the value corresponding to the WebbPSF [^5] https://www.stsci.edu/jwst/science-planning/
@@
F [^4] W FWHM, 0.14”) and the sixth one (corresponding to the larger aperture in the UNCOVER catalog by Weaver et al., 2024). We point out that we did not correct the fluxes for the cluster magnification effect in the ABELL2744 field.
Figure 4: Number counts in four NIRCam bands (total magnitudes from the optimal catalogs, normalized to field area); black thick line is the total of all fields.
For each field, we provide an “optimal” catalog, in which the total flux of each source is obtained from the colors computed in a preferred aperture chosen on the basis of the object segmentation area. We used SExtractor ISOAREA_IMAGE as a proxy for it, and selected as the preferred aperture the one immediately larger than the value $\sqrt{\texttt{ISOAREA\_IMAGE}/\pi}$ (this implies that the diameter of the optimal aperture is close to the radius of the circularized detected area). The value of this preferred aperture is also reported in the catalog.
@@ 3.4.1 PSF models
Figure 5: PSF “über”-models for all the bands in the catalogs. See text for details.
A major difference with respect to our previous efforts is the usage of empirical PSFs rather than models simulated using the WebbPSF application. For each band, we created empirical models stacking isolated, high SNR stars, singled out visually after an automatic pre-selection performed using catalogs created with ad-hoc SExtactor runs on each band. After trying various options, among which using stars from each field to create the PSFs for that field only, we found that the best results in terms of growth curves and photometric vs. spectroscopic redshifts comparison (see Section 5) was to create an “über”-model for each band, stacking all the available good stars from all fields after rotating each one by the average position angle (PA) of the observation, and finally rotating back the obtained model to the PA of the considered field (we used the Python module numpy.rotate for this task). In the cases in which the observations are stacks of different epochs, we created models using the stars with the most common orientation in our selection (also considering those consisting of two superposed PSFs with different orientations), which typically sample the largest area in the field. Fig. 5 show the final PSF “über”-models before the final rotation for being used in the different fields.
@@ 3.4.2 Background subtraction
a-phot can perform “on-the-fly” local background subtraction while measuring fluxes. Including this feature is not necessarily the best choice in all cases, as it may yield sub-optimal estimates close to bright sources or in densely populated regions; however, it is typically reasonable to apply it. Therefore, we release both sets of catalogs with the local background subtraction option switched on and off, but we suggest to use the background-subtracted catalogs as best choice. We checked that the difference is not dramatic for the vast majority of the sources, but is has an impact in some cases. Fig. 6 shows the difference in measured fluxes in six bands with and without a-phot background subtraction, in all fields: major differences can be seen around the bright galaxies in the ABELL2744 cluster (notice how the a-phot background subtraction makes the objects closest to the cluster members fainter, but those slightly farther away brighter, compensating for over-subtraction in the image processing phase) and in an extended area of the PRIMER-UDS field.
@@ 3.4.3 Galactic extinction
We finally corrected all the total fluxes for the effects of galactic extinction, taking advantage of the calculator provided by the NASA/IPAC Extragalactic Database (NED) [^6] https://ned.ipac.caltech.edu/extinction_calculator, which gives the average dimming in each band at any given equatorial coordinates (we provide those corresponding to the center of the FoVs of the fields). We use such values to compute the extinction-corrected fluxes, and use the latter for the final catalogs.
Fig. 4 shows the number counts of all fields in four bands using the magnitudes obtained using the total fluxes in the “optimal” catalog and after the correction for galactic extinction. Comparing it with Fig. 3, it can be seen that the counts in the F [^4] W and F356 bands typically peak close to the 50% completeness detection magnitude, which is consistent with previous studies (see e.g. Guo et al., 2013, their Fig. 4).
Table 5: Flags assigned in the catalogs. The final flag is the sum of the individual flagging values.
@@
+1-8 Source is missing HST bands
+10-80 Source is missing JWST bands
+100 Source is contaminated by close neighbors, or has bad pixels in detection
@@
3.5 Star/galaxy separation, spurious detections, and flagging
@@
To identify and flag point-like sources and potentially spurious detections (PLS and PSD, respectively) we adopted a Principal Component Analysis (PCA) technique, described in details in Appendix C. In short, taking advantage of the quantities MU_MAX, MAG_AUTO, MAGERR_AUTO and FLUX_RADIUS estimated with SExtractor on the F356W+F [^4] W detection stack, we obtained a 2-dimensional projection space in which PLS occupy an extremely well-defined and tight locus, and PSD gather in a relatively confined region. The first column of Fig. 9 shows the resulting diagram for all the fields, with the inner subplot magnifying the region where spurious and regular sources overlap; the exact formulae used to isolate PLS and PSD in the diagram are given in Sect. C. The second and third columns show the position of the sources in two complementary diagnostic planes, i.e. the MU_MAX - MAG_AUTO plane and the FLUX_RADIUS vs SNR plane, where PLS and PSD also occupy well-defined regions; the usage of the PCA approach allows for a more general identification, combining the two in a fluid way.
@@
We then assigned a further flag to each source, based on the detection and photometric measurements. Table 5 describes the used values, the final flag being the sum of all the individual addenda. Power-of-two values multiplied by 100 are assigned on the basis of the detection measurements, as output by the a-phot code, the maximum value being 1500; while addenda below 100 indicate the number of HST and/or JWST bands missing because of the different observational coverage of the areas. We also included a special flag for the sources with all HST measurement having negative values (a few sources in a limited region of the ABELL2744 field, because of problematic background subtraction in the cluster core area). To give an example, a source flagged with the value 11016 identifies a point-like object, which is blended and saturated in the detection band, and has one JWST and six HST bands missing. Thus, a flag lower than 199 typically indicates a “regular” galaxy-like source which might have missing bands and/or be contaminated in its detection total flux, but it is not blended in detection, not saturated, not close to the borders of the image, and not identified as a star or a spurious detection. Clearly, these flags are a useful diagnostic tool, but we suggest to use them with caution, as they are the result of many automatic processes that cannot be 100% accurate.
A certain number of spurious detections is unavoidably destined to remain in the catalogs. To further improve the purity, we visually checked all sources with photometric redshift estimate above 10 (see Sect. 5), to exclude at least the most obvious errors in this important sub-space of the catalogs. Doing so, we found that most remaining cases were defects in the detection stack, either caused by missing coverage in the F356W or F [^4] W images, and/or by obvious reduction errors or noise features that were not singled out in the diagnostic planes. The most affected fields were ABELL2744, CEERS and PRIMER-UDS.
@@
It is reasonable to assume that similar spurious detections could exist at photo-$z<$10, although detecting in the reddest bands favours high photometric redshift estimates for these kind of fake sources. However, since it would be impossible to go through all of the catalogs to find them, we invite the users to carefully inspect any objects selected for scientific purposes.
4 Validation of photometry

### v3

@@ 2.4 PRIMER-COSMOS and PRIMER-UDS
Data is from the PRIMER programme (GO 1837, P.I. Dunlop; no F775W for PRIMER-COSMOS and no F775W and F105W for PRIMER-UDS). The COSMOS FoV has an area of $\simeq$141.8 sq. arcmin, and the UDS FoV has an area of $\simeq$251.2 sq. arcmin. The images have been reduced by D. Magee, with ancillary HST data re-reduced from the CANDELS imaging, UVCANDELS (Wang et al., 2024b), and supplemental programme 16872 (P.I. Grogin).
@@ 3 Methods
In this section, we summarise the techniques and algorithms we used to prepare the images and to extract photometric information.
@@ 3.1 Alignment and astrometry
We assessed the quality of the astrometric registration by cross-matching the source coordinates ALPHAWIN_J2000 and DELTAWIN_J2000 obtained by running SExtractor on all the bands in each field before re-projecting the images to the same common grid. In most cases we found small $\Delta$RA and $\Delta$DEC offsets (of the order of a few mas) between the native astrometry of the JWST mosaics (which are aligned to Gaia) and the HST bands. However, since in some cases the offsets were larger and not negligible, we applied a custom algorithm to correct all of them. First, we computed the offset between the HST WFC3 F160W band and the JWST NIRCam F150W band, and re-aligned the F160W image (we applied a rigid shift to the whole image). Then we corrected the offsets of the other HST bands by cross-matching the coordinates of the sources with those extracted from the re-aligned F160W band (see Fig. 1 and Table B.1). Finally, we re-projected all the images on the same NIRCam pixel grid.
Figure 1: Example of the validation tests made to fix the astrometric registration of the HST bands. Shown is the displacement $\Delta$RA and $\Delta$DEC between the sources detected in the CEERS F125W image and those detected in the CANDELS EGS F160W image, re-aligned to the reference frame of the CEERS F150W image, before (left panel) and after (right panel) applying the correction described in Sect. 3.1.
@@ 3.2 RMS scaling
Because the photometric errors were computed by means of the RMS maps (see Sect. 3.4), we checked that the latter were indeed representative of the true uncertainties of the measurement, using an improved version of the technique described in M22. In short, the RMS image of each band was subdivided into two to four complementary sub-regions of comparable exposure time, by means of an automatic algorithm applied to the weight maps. Then, in each of these sub-region, 300 artificial point sources (WebbPSF simulated point spread functions) were injected at random positions, excluding areas assigned to real sources by means of SExtractor segmentation maps. The dispersion of their fluxes (measured within an aperture of 0.1” using the software a-phot, Merlin et al., 2019) was compared to their nominal errors to obtain a re-scaling factor for that region of the RMS map. Finally, the original map (which includes Poissonian photon noise) was then re-scaled by the median value of such factors. The typical values for SW bands are below $\sim$1.5, while they can get as high as $\sim$2.0 for some LW bands and fields.
[…]
@@ 3.3 Detection
Because the catalogues have mainly been designed for the study of the high-redshift Universe, major effort was put into optimising the measurements of faint extended objects, rather than those of bright local galaxies or stars. The strategy is similar to that adopted in M22 and P23: we chose to carry out our detection in the infrared, smoothing the scientific image with a Gaussian convolution filter with full width at half maximum (FWHM) of 0.14” (close to the F356W and F [^4] W FWHMs; see below), and applying a detection threshold corresponding to S/Nsimilar-$\sim$2, to pick up as many faint high-redshift sources as possible. However, there is a substantial difference with respect to previous work: rather than just using F [^4] W as the detection band, we created a weighted stack of F356W and F [^4] W, allowing us to single out sources that peak at 3.5 to 4 $\mu$m, and exploiting the F356W mosaics, which are often as deep as (or deeper than) the F [^4] W ones (see Fig. 2). The resulting depths of the detection stacks are shown in Fig. 3.
@@
We then ran SExtractor v2.8.6, in the customised version used for the CANDELS campaign. Most of the parameters were left at the default values, with the exception of those given in Table 3. We tried to use the same values for all fields, but we found that DETECT_THRESH and ANALYSIS_THRESH needed adjustment in some cases, to ensure optimal results. In particular, we set them to 0.85 for PRIMER-COSMOS, and to 0.55 for ABELL2744; these values yielded the best trade-off between purity and completeness, allowing for the detection of faint sources, while avoiding too many spurious objects from being included on the list.
@@
The final detection catalogues contain a grand total of 531,173 sources, as reported in Table 4. In the catalogues we include a unique object identifier number, the equatorial position (right ascension and declination, in degrees), and basic morphological information. The latter includes the area of the segmented cluster of pixels and the half light radius from SExtractor, while the semi-major axis, ellipticity, and position angle of the elliptical isophote were obtained by means of a-phot (see Appendix A for more details).
@@
Figure 4 shows the point-source detection completeness for the six fields, determined by injecting fake PSF-shaped objects in empty regions of the F356W+F [^4] W stack (using the detection segmentation map as a mask), running SExtractor with the same parameters used in the actual detection process, and checking the fraction of them being actually detected. The different depths of the detection images is evident, with NGDEEP being the deepest with 90% at ABsimilar-$\simeq$30 and 50% at ABsimilar-$\simeq$30.5 and PRIMER-UDS the shallowest, with 90% at ABsimilar-$\simeq$28.5 and $50$% at ABsimilar-$\simeq$29. We note the complicated pattern of the JADES-GS field, which comprises deep and shallow regions (90% completeness at ABsimilar-$\simeq$28.5, but 50% at ABsimilar-$\simeq$30).
@@
Figure 4: Detection completeness: Fraction of fake point sources of a given magnitude injected in the detection F356W+F [^4] W stacks and detected with SExtractor runs performed with the parameters used in real runs.
@@ 3.4 Photometry
As in our previous efforts, in our public catalogues we provide total fluxes. The main methods used for this work are the same ones adopted in M22 and P23. In short, we compute colours in fixed circular apertures on PSF-matched images; as in P23, we smoothed all bands to the F [^4] W resolution, except for the HST WFC3 bands which, having a broader FWHM, were smoothed to the F160W resolution. Assuming no colour gradients outside the apertures, we then scale to total fluxes, multiplying the colour term by the flux within Kron (1980) elliptical apertures in the F356W+F [^4] W stack as measured by the software a-phot (Merlin et al., 2019). So, the total flux in each band is given by the formula $f_{tot,band}=c_{ap,det}\times f_{ap,band}$, where $c_{ap,det}=f_{tot,det}/f_{ap,det}$. As pointed out in M22, our a-phot Kron-like aperture on average tends to gather more light than the standard SExtractor MAG_AUTO (see Fig. 4 in M22), so we did not apply any further aperture correction to the total fluxes. Errors are estimated by summing in quadrature the relevant pixels from the RMS maps and applying the same formula, $e_{tot,band}=c_{ap,det}\times e_{ap,band}$; this choice is motivated by the fact that most scientific applications (e.g. SED-fitting or colour-selections) are essentially based on colours, so the propagation of the total flux error would overestimate the relevant uncertainty. We computed the fluxes within nine apertures (with diameters 0.2”, 0.28”, 0.33”, 0.50”, 0.66”, 0.70”, 1.32”, 2.65”, and 5.30”), and built just as many catalogues of total fluxes. The values of the apertures correspond to integer multiples of the F [^4] W FWHM 0.165”, except for the two smallest ones (a fixed 0.1” radius aperture and the value corresponding to the WebbPSF [^9] https://www.stsci.edu/jwst/science-planning/
@@
F [^4] W FWHM, 0.14”) and the sixth one (corresponding to the larger aperture in the UNCOVER catalogue by Weaver et al., 2024). We point out that we did not correct the fluxes for the cluster magnification effect in the ABELL2744 field.
For each field, we provide an ‘optimal’ catalogue, in which the total flux of each source is obtained from the colours computed in a preferred aperture chosen on the basis of the object segmentation area. We used SExtractor ISOAREA_IMAGE as a proxy for it, and selected as the preferred aperture the one immediately larger than the value $\sqrt{\texttt{ISOAREA\_IMAGE}/\pi}$ (this implies that the diameter of the optimal aperture is close to the radius of the circularized detected area). The value of this preferred aperture is also reported in the catalogue.
@@ 3.4.1 PSF models
Figure 5: PSF “über”-models for all the bands in the catalogues. See text for details.
A major difference with respect to our previous efforts is the usage of empirical PSFs, rather than models simulated using the WebbPSF application. For each band, we created empirical models stacking isolated, high-S/N stars, singled out visually after an automatic pre-selection performed using catalogues created with ad-hoc SExtactor runs on each band. After trying various options, including using stars from each field to create the PSFs for that field only, we found that the best results in terms of growth curves and photometric versus spectroscopic redshifts comparison (see Sect. 5) was to create an ‘über’ model for each band, stacking all the available good stars from all fields after rotating each one by the average position angle (PA) of the observation and, finally, rotating back the obtained model to the PA of the considered field (we used the Python module numpy.rotate for this task). In cases where the mosaics are made up of stacks of images from different epochs, we created models using the stars with the most common orientation in our selection (also considering those consisting of two superposed PSFs with different orientations). These models typically sample the largest area in the field. Figure 5 shows the final PSF ‘über’ models before the final rotations for their usage in the different fields.
@@ 3.4.2 Background subtraction
a-phot can perform ‘on-the-fly’ local background subtraction, while measuring the fluxes. Including this feature is not necessarily the best choice in all cases, as it may yield sub-optimal estimates close to bright sources or in densely populated regions; however, it is typically reasonable to apply it. Therefore, we released both sets of catalogues with the local background subtraction option switched on and off, but we suggest using the background-subtracted catalogues as the optimal choice. We checked that the difference is not dramatic for the vast majority of the sources, but it does have an impact in some cases. Figure 6 shows the difference in measured fluxes in six bands with and without a-phot background subtraction, in all fields. Major differences can be seen around the bright galaxies in the ABELL2744 cluster; we note how the a-phot background subtraction makes the objects closest to the cluster members fainter, but those slightly farther away brighter, compensating for over-subtraction in the image processing phase. Such differences are also seen in an extended area of the PRIMER-UDS field.
@@ 3.4.3 Galactic extinction
Finally, we corrected all the total fluxes for the effects of galactic extinction, taking advantage of the calculator provided by the NASA/IPAC Extragalactic Database (NED) [^10] https://ned.ipac.caltech.edu/extinction_calculator, which gives the average dimming in a number of bands, at any given equatorial coordinates (we provide those corresponding to the centre of the FoVs of the fields). NIRCam filters are not included in the list, so for them, we interpolated between the available bands at the closest wavelengths. We used such values to compute the extinction-corrected fluxes and included the latter for the final catalogues.
Figure 7: Number counts in four NIRCam bands (total magnitudes from the optimal catalogues, normalised to field area). The thick black line is the total of all fields.
3.4.4 Number counts
Figure 7 shows the number counts of all fields in four bands using the magnitudes obtained using the total fluxes in the ‘optimal’ catalogue and after the correction for galactic extinction. Comparing it with Fig. 4, it can be seen that the counts in the F [^4] W and F356 bands typically peak close to the 50% completeness detection magnitude, which is consistent with previous studies (see e.g. Guo et al., 2013, their Fig. 4).
Table 5: Flags assigned in the catalogues.
@@ 3.4.4 Number counts
+1-8 Source is missing HST coverage
+10-80 Source is missing JWST coverage
+100 Source is contaminated by close neighbours, or has bad pixels in detection
@@
 [^11] The final flag assigned to each source is the sum of the individual flagging values.
3.5 Point-like sources, spurious detections, and flagging
Figure 8: Example of the PCA diagram used for the identification of point-like (red points) and spurious (black points) sources.
@@ 3.5 Point-like sources, spurious detections, and flagging
To identify and flag point-like sources and potentially spurious detections (PLS and PSD, respectively) we adopted a principal component analysis (PCA) technique, described in detail in Appendix C. In short, taking advantage of the quantities MU_MAX, MAG_AUTO, MAGERR_AUTO and FLUX_RADIUS estimated with SExtractor on the F356W+F [^4] W detection stack, we obtained a two-dimensional projection space in which PLSs occupy an extremely well-defined and tight locus and PSDs gather in a relatively confined region. The first column of Fig. C.1 shows the resulting diagram for all the fields, with the inner subplot magnifying the region where spurious and regular sources overlap; an example for one field (PRIMER-UDS) is given in Fig. 8. The exact formulae used to isolate PLS and PSD in the diagram are given in Appendix C and Table C.1. The second and third columns of Fig. C.1 show the position of the sources in two complementary diagnostic planes, namely: the MU_MAX - MAG_AUTO plane and the FLUX_RADIUS versus S/N plane, where PLS and PSD also occupy well-defined regions. The application of the PCA approach allows for a more general identification, combining the two in a fluid way.
@@
We then assigned a further flag to each source, based on the detection and photometric measurements. Table 5 describes the used values, the final flag being the sum of all the individual addenda. Power-of-two values multiplied by 100 were assigned on the basis of the detection measurements, as an output by the a-phot code, the maximum total value being 1500. Addenda below 100 indicate the number of HST and/or JWST bands missing because of the different observational coverage of the areas. We also included a special flag for the sources with all HST measurement having negative values (a few sources in a limited region of the ABELL2744 field, because of problematic background subtraction in the cluster core area). To give an example, a source flagged with the value 11016 identifies a point-like object, which is blended and saturated in the detection band, and has one JWST and six HST bands missing. Thus, a flag lower than 199 typically indicates a ‘regular’ galaxy-like source which might have missing bands and/or be contaminated in its detection total flux, but it is not blended in detection, not saturated, not close to the borders of the image, nor identified as a star or a spurious detection. Clearly, these flags are a useful diagnostic tool, but we suggest using them with caution, as they are the result of many automatic processes that cannot achieve full accuracy. A certain number of spurious detections is unavoidably destined to remain in the catalogues. To further improve the purity, we visually checked all sources with photometric redshifts estimate above 10 (see Sect. 5) to exclude at least the most obvious errors in this important sub-space of the catalogues. In doing so, we found that most remaining cases were defects in the detection stack, either caused by missing coverage in the F356W or F [^4] W images and/or by obvious reduction errors or noise features that had not been singled out in the diagnostic planes. The most affected fields were ABELL2744, CEERS, and PRIMER-UDS.
@@
It is reasonable to assume that similar spurious detections could exist at photo-$z$$<$10, although making detections in the reddest bands favours high photometric redshift estimates for these kinds of fake sources. However, since it would be impossible to go through all of the catalogues to find them, we invite users to carefully inspect any objects selected for scientific purposes.
4 Photometry validation